If u r intelligent ans dis
When I was 2, my brother was half my age, now I am 100, how old is he?

Answer by checkley79(3341)

YOU ARE 2 YEARS OLD.
YOUR BROTHER IS 2(1/2)=1 YEAR OLD.
IN 98 YEARS YOU ARE 2+98=100 YEARS OLD.
YOUR BROTHER IN 98 YEARS IS 1+98=99 YEARS OLD.